# Hierarchical Graph Neural Network for Compressed Speech Steganalysis

### v1

@@ Abstract
Steganalysis methods based on deep learning (DL) often struggle with computational complexity and challenges in generalizing across different datasets. Incorporating a graph neural network (GNN) into steganalysis schemes enables the leveraging of relational data for improved detection accuracy and adaptability. This paper presents the first application of a Graph Neural Network (GNN), specifically the GraphSAGE architecture, for steganalysis of compressed voice-over-IP (VoIP) speech streams. The method involves straightforward graph construction from VoIP streams and employs GraphSAGE to capture hierarchical steganalysis information, including both fine-grained details and high-level patterns, thereby achieving high detection accuracy. Experimental results demonstrate that the developed approach performs well in uncovering quantization index modulation (QIM)-based steganographic patterns in VoIP signals. It achieves detection accuracy exceeding 98% even for short 0.5-second samples, and 95.17% accuracy under challenging conditions with low embedding rates, representing an improvement of 2.8% over the best-performing state-of-the-art methods. Furthermore, the model exhibits superior efficiency, with an average detection time as low as 0.016 seconds for 0.5-second samples—an improvement of 0.003 seconds. This makes it efficient for online steganalysis tasks, providing a superior balance between detection accuracy and efficiency under the constraint of short samples with low embedding rates.
@@
[label1]organization=LCPTS Laboratory, University of Sciences and Technology Houari Boumediene (USTHB),addressline=P.O. Box 32, city=El-Alia, Bab-Ezzouar, postcode=16111, state=Algiers, country=Algeria
@@
[label2]organization=LSEA Laboratory, Department of Electrical Engineering, University of Medea,postcode=26000, country=Algeria
@@
[label3]organization=Cybersecurity Institute, School of Computer Science and Informatics, University of Liverpool,postcode=L69 3BX, country=UK
@@ 1 Introduction
Data hiding has garnered significant attention from researchers in recent years, spanning various disciplines. On the one hand, it serves to verify integrity and ensure information remains unaltered during transmission, as seen in watermarking [hemis2018adjustable, rebahi2023image]. On the other hand, it can be utilized to conceal confidential information through unconventional means, exemplified by steganography [kheddar2022high, abdelmged2016improving]. Steganography and steganalysis are two complementary aspects of covert communication. While steganography involves concealing secret information within apparently benign carriers such as images [megias2022subsequent], text [khosravi2019new], and speech [kheddar2022high], steganalysis is focused on detecting and unveiling such concealed communication [kheddar2024deep]. Over the years, the proliferation of Internet services has facilitated the transmission of multimedia data over digital networks. Among these services, voice over internet protocol (VoIP) has emerged as a prominent means of enabling real-time voice communication over the Internet. This increased reliance on VoIP, driven by the success of platforms like Skype, WhatsApp, and Zoom, has made it an attractive choice for concealing and transmitting hidden information [tian2014improving]. The ubiquity and high-volume nature of VoIP traffic make it an appealing vector for steganographic purposes, presenting both opportunities for clandestine communication and challenges for cybersecurity professionals.
@@
In VoIP communications, speech is typically compressed using codecs, such as G.722, G.723, and G.729, to reduce bandwidth requirements. This compression process, which involves the quantization of speech parameters, inadvertently creates potential vulnerabilities that can be exploited for steganographic purposes. Malicious actors, including cybercriminals and extremist groups, may exploit these vulnerabilities by manipulating VoIP software to facilitate covert communications. This scenario presents a significant challenge to communication monitoring and network security. Steganography for VoIP can occur both in network protocol fields and payload data [wu2021steganography, kheddar2019pitch]. However, while the former approach may provide some degree of concealment, it is the latter that takes precedence [ghanem2022towards, farzaan2025ai, ghanem2015towards, ghanem2016enhancing, basnet2025advanced]. This distinction arises from the fact that the practice of concealing information within network protocol fields, which encompass many layers of the open systems interconnection (OSI) model, ultimately provides less robust concealment capabilities. These protocol fields are typically public, and their data tends to remain fixed under most ordinary circumstances, rendering them less secure for covert communication [yang2019hierarchical]. On the contrary, the method of embedding data within payload data, characterized by its dynamic nature and temporal variations, offers a substantially higher degree of concealment. This dynamism makes it considerably more challenging to detect covert practices within VoIP streams, reinforcing its position as the preferred choice for steganographic endeavors [yang2019hierarchical, yang2022real]. Additionally, as steganography methods become more sophisticated, particularly with the advent of AI-generated content, advanced steganalysis techniques are necessary to keep pace [li2022general].
@@
An effective steganalysis in the context of VoIP streams should meet two crucial requirements [lin2018rnn, ghane2023novel]. Firstly, it should operate in real-time, ensuring that the time required for detection is minimized due to the need for swift action against potential malicious activities. Secondly, the steganalysis method must be capable of detecting short samples of VoIP streams, as covert information may be hidden within brief segments of the communication. Adding to these two requirements, it should be sensitive enough to detect low embedding rates, as some covert data may be hidden with minimal changes to the host signal. Meeting these stringent requirements represents a significant and persistent challenge in developing practical and reliable VoIP steganalysis methods, which is essential to ensure their successful integration into real-world applications [lin2018rnn, wu2021steganography]. In this work, we address the core research problem of designing a steganalysis approach for VoIP that can achieve real-time detection performance, maintain high accuracy on short speech segments, and remain effective under low embedding rates.
@@ 1.1 Motivation
With the rapid advancement of deep learning (DL), recent methods have leveraged deep models to automatically learn discriminative features. Various neural network architectures, primarily leveraging convolutional neural network (CNN)- and recurrent neural network (RNN)-based designs [qiu2022steganalysis, lin2018rnn, yang2022real, wu2023mfpd], have been explored in VoIP steganalysis. These models are specifically designed to capture distinct codeword correlation features between cover and stego VoIP streams, achieving state-of-the-art (SOTA) detection outcomes. While effective in capturing sequential (RNNs) or local spatial (CNNs) patterns, these traditional deep learning architectures face inherent limitations when dealing with the relational structure of compressed VoIP data affected by steganography. QIM steganography, in particular, subtly alters the dependencies and correlations between line spectral frequency (LSF) codewords across frames and within their neighborhoods. CNNs, designed for grid-like data, struggle to explicitly model these non-Euclidean, inter-frame relationships. RNNs, while adept at sequences, might not efficiently capture the graph-like dependencies that extend beyond simple linear order or local windows, especially when the crucial information lies in the way codewords relate to each other rather than just their individual values or linear sequence.
Recently, significant attention has been directed toward adapting DL methods for graph-structured data, giving rise to the the emergence of graph neural networks (GNNs) as a prominent topic [wu2020comprehensive]. GNNs are uniquely suited for modeling complex relationships within non-Euclidean data by explicitly learning from graph structures, where nodes represent entities (e.g., LSF codeword frames) and edges represent their relationships (e.g., temporal dependencies). This intrinsic capability allows GNNs to capture both fine-grained local dependencies and high-level global patterns by aggregating information from connected neighborhoods. GNNs have demonstrated efficiency in representing and analyzing graph data across various computer vision fields, including action recognition [geng2024hierarchical], object tracking [zhan2024gnn], and natural language processing, such as text classification [peng2024novel]. However, there have been limited efforts in applying GNNs to steganalysis, primarily in the domains of images, as in [liu2022graph, liu2023jpeg], and text, as in [wu2021linguistic, fu2022hga]. This highlights a significant gap where GNNs’ relational modeling power could be particularly advantageous for VoIP steganalysis.
@@ 1.2 Contribution and paper structure
    It achieves detection accuracy exceeding 98% even for short 0.5-second samples, and 95.17% accuracy under challenging conditions with low embedding rates, representing an improvement of 2.8% over the best-performing state-of-the-art methods. Furthermore, the model exhibits superior efficiency, with an average detection time as low as 0.016 seconds for 0.5-second samples—an improvement of 0.003 seconds.
@@ 2 Related work
Vocoders, such as G.723 and G.729, are widely applied in VoIP communication with the aim of lowering decoding errors via the analysis-by-synthesis (AbS) approach, maintaining excellent speech quality while achieving high compression ratios. In contrast to conventional steganographic methods typically used for images or text, low bit-rate (LBR) VoIP streams, characterized by the AbS linear predictive coding technique, pose a unique challenge due to their minimal redundancy within the encoded speech. To address this, researchers have explored two distinct categories of data hiding techniques, each linked to different stages in the encoding process. The first category involves altering specific elements in the compressed speech stream, often utilizing methods like least significant bit (LSB) replacement, such as in [kheddar2019pitch]. However, this is complicated due to the limited data-hiding space in LBR speech. The second category focuses on embedding covert information during encoding, by altering encoding features, notably the linear predictive coding (LPC) [laskar2024enhancing], fixed codebook (FCB) [liu2016neighbor], and adaptive codebook (ACB) [shufan2015steganography].
@@
In the second category, QIM-based steganography is most commonly employed, especially within LPC domains. QIM steganography, originally developed by Chen et al. [chen2001quantization], is widely used in literature. It conceals steganograms by altering the quantization vector within the speech code, introducing minimal unperceptible distortion and offering high data hiding capabilities. Thus, when it comes to embedding information in VoIP streams, QIM emerges as a suitable scheme that poses a significant challenge for detection.
@@
On the one hand, research into QIM steganography detection methods has primarily concentrated on image carriers [malik2011nonparametric]. Subsequently, a limited number of research studies have been proposed in the field of steganalysis, particularly focusing on QIM-based steganography in VoIP, such as in [yang2019hierarchical, sun2021audio]. Due to the distinct characteristics of VoIP streams, the detection methods proposed for use in images cannot be directly applied within the VoIP context. Moreover, classical audio steganalysis methods based on statistical feature extraction in the uncompressed domain are inapplicable to VoIP. This is due to the minimal additional distortion introduced by VoIP QIM steganography in decoded speech signals, making it challenging to extract suitable features for steganalysis [yang2019steganalysis]. To address this issue, some research in the VoIP field has proposed the use of handcrafted statistical features to detect significant changes in codewords used in LBR speech codec streams caused by QIM steganography. For instance, Li et al. [li2012steganalysis] examined the impact of QIM steganography on G.729, where it altered the quantization index of the LPC filter. They employed a statistical model to capture codeword distribution characteristics and, in conjunction with a support vector machine (SVM), formulated an efficient system for steganographic detection. In a related study, Li et al. [li2012detection] developed the index distribution characteristics (IDC) steganalysis method, employing codeword distribution histograms and first-order Markov chain-derived state transition probabilities as correlation features. These features, integrated with SVM, were utilized to classify steganographic samples with a focus on inter-frame transition probabilities. Expanding upon this concept, the studies by Li et al. [li2017steganalysis] and Wu et al. [wu2020steganalysis] aim to detect steganograms embedded using QIM techniques. Both works employ principal component analysis (PCA) for reducing feature dimensionality and SVM for classification. Notably, Li et al. [li2017steganalysis] focus on state transition probabilities of intra-frame codewords as features, while Wu et al. [wu2020steganalysis] suggested extracting features from the original speech signal by analyzing its distribution patterns and transmission likelihoods. Meanwhile, Yang et al. [yang2018steganalysis] employed temporal dependencies between codewords across frames to model a Codeword Bayesian Network (CBN), where the parameter learning process was guided by a Dirichlet prior. Yang et al. [yang2019fast] proposed a fast steganalysis which involved mapping vector quantization codewords to a semantic space, followed by feature extraction using a hidden layer input into a softmax classifier. Such methods require meticulous feature design to align with the characteristics of speech data and specific information hiding techniques. Consequently, they struggle to effectively address generative steganography that produces high-quality stego speech. On the other hand, the proposed GraphSAGE-based algorithm addresses these challenges by utilizing relational data through GNNs and efficiently managing the minimal distortion and short sample lengths typical of VoIP streams.
@@
Besides, recent advancements in DL have led to improved robustness and effectiveness in steganalysis methods, which can be broadly categorized into two main categories: general steganalysis approaches and specific steganalysis approaches. In the first category, several methods have been introduced to detect a range of steganographic techniques. For example, Hu et al. [hu2021detection] addressed heterogeneous parallel steganography (HPS) by introducing the steganalysis feature fusion network (SFFN). The architecture of SFFN comprises three core modules: a network for feature extraction, a component for feature integration, and a classification unit, effectively extracting and combining steganalysis features from HPS methods for dependable predictions. Building upon this, Wang et al. [wang2021fast] proposed a fast and efficient steganalysis method which leveraged attention mechanisms. Their suggested method adeptly extracts key characteristics related to steganography exceptions and fuses these features targeting multiple steganography techniques utilized in HPS. Li et al. [li2021detection] introduced a unified detection framework, termed CBCA, which leverages codeword embedding along with bidirectional long short-term memory (Bi-LSTM) and CNN-based attention mechanisms. This approach is capable of identifying three distinct steganographic techniques concurrently. Wang et al. [wang2022steganalysis] proposed a steganalysis method aimed at identifying various steganographic schemes within compressed speech signals. Their framework introduces a codeword-distributed embedding component designed to generate compact representations from compressed codewords. To capture dependencies across different contexts, it employs two correlation extraction modules: a global-guided unit that incorporates Bi-LSTM and multi-head self-attention, and a local-guided unit built using the convolutional block Attention module (CBAM), enabling the modeling of both global and local correlations before and after data hiding. Detection is accomplished through fully connected layers, ensuring accurate identification. Li et al. [li2022general] introduced a generic frame-level steganalysis approach for LBR steganography, utilizing dual-domain feature representation and a Transformer-based model named Stegaformer. The proposed method consists of two modules, dual-domain representation and Stegaformer. Li et al. [li2023sanet] introduced SANet, an independent steganalysis network for speech encoding and steganography. This approach unifies compressed speech outputs from various codecs by converting them into a common uncompressed-domain format, introducing an intermediate representation. A Bi-LSTM neural network is developed to extract steganography-sensitive characteristics through collaborative correlation features. This approach achieves SOTA detection performance for various steganography algorithms across different speech encoders. Recently, Wang et al. [wang2025swan] proposed E-SWAN, a deep learning-based sliding window analysis network that combines LSTM and convolutional modules for real-time VoIP speech steganalysis. The model demonstrates competitive performance in detecting two steganography methods. Similarly, Lin et al. [li2025compressed] introduced a steganalysis approach that integrates Bi-LSTM, 3D convolution and attention mechanisms to detect two QIM-based steganography variants in VoIP streams. Although these algorithms offer some capability for broad steganalysis applications, their performance often falters when confronted with specific steganographic techniques, particularly QIM. In such specialized scenarios, these general methods typically exhibit lower accuracy, highlighting the need for more targeted approaches.
@@
Consequently, several researchers prefer to exploit the power of DL to design specialized steganalysis algorithms, particularly geared toward QIM-based steganography. For example, Lin et al. [lin2018rnn] identified four robust codeword correlation patterns and introduced a steganalysis model based on RNN for improved detection accuracy. Nonetheless, their model, RNN-SM, centred on global contextual information, utilizing a two-layer LSTM. Yang et al. [yang2019steganalysis] proposed a novel approach to enhance local representations that combined RNN and CNN, resulting in the CNN-LSTM model, which outperformed RNN-SM in the detection of QIM-based steganography. Yang et al. [yang2019hierarchical] incorporated an attention mechanism within a hierarchical convolutional structure, further enhancing steganalysis results. In a subsequent study [yang2020fcem], the same authors presented a lightweight neural architecture called the fast correlation extract model (FCEM), which integrates positional encoding and multi-head attention to capture correlation features. Qiu et al. [qiu2022steganalysis] developed an efficient steganalysis that incorporates a codeword embedding layer to capture dense representations, employs a bidirectional LSTM layer and incorporates a gated attention module to model contextual feature distributions. Yang et al. [yang2022real] designed a multi-channel convolutional sliding window mechanism to capture inter-frame dependencies between a target frame and its adjacent frames. Recently, the utilization of a transformer encoder [zhang2023tenet] and federated learning [tian2023fedspy] has been implemented for the detection of QIM-based steganography in G.729 speech encoders. These two approaches demonstrated SOTA detection performance. Zhang et al. [zhang2024spm] tackled the challenge of identifying payload locations in QIM-based VoIP steganography by proposing an LSTM-based steganalysis method named SPM.
@@
The mentioned studies underscore the adoption of DL-based methods as the benchmark for real-time VoIP steganalysis, demonstrating superior detection capabilities and efficiency in comparison to handcrafted features. The integration of DL with graph data has led to the widespread adoption of GNNs. These DL-based models, specialized for graph domains, are crafted to capture representations of graph data adeptly, proficiently handling tasks centred on both individual nodes and entire graphs. GNNs capture relationships within graphs by exchanging messages among nodes and maintaining a state that encodes information from any connected neighborhood through edges [zhou2020graph]. Furthermore, GNNs exhibit an enhanced capability to leverage global information, contributing to their compelling performance. While GNNs have been recently employed in various studies for image [liu2022graph, liu2023jpeg] and text steganalysis [wu2021linguistic, fu2022hga, pang2023cats], there is a noticeable gap in research focusing on steganalysis in the context of VoIP.

### v2

@@ Abstract
Steganalysis methods based on deep learning (DL) often struggle with computational complexity and challenges in generalizing across different datasets. In the specific case of voice-over-IP (VoIP) speech streams, detection is particularly challenging because the low bit-rate encoding creates complex, relational dependencies between speech frames. Conventional DL models, which treat data as simple sequences or grids, often fail to capture these complex inter-frame dependencies effectively. To address this gap, this paper presents the first application of a graph neural network (GNN), specifically the GraphSAGE architecture, for steganalysis of compressed VoIP speech streams. The method involves straightforward graph construction from VoIP streams and employs GraphSAGE to capture hierarchical steganalysis information, including both fine-grained details and high-level patterns, thereby achieving high detection accuracy. Experimental results demonstrate that the developed approach performs well in uncovering quantization index modulation (QIM)-based steganographic patterns in VoIP signals. It achieves detection accuracy exceeding 98% even for short 0.5-second samples, and 95.17% accuracy under challenging conditions with low embedding rates, representing an improvement of 2.8% over the best-performing state-of-the-art methods. Furthermore, the model exhibits superior efficiency, with an average detection time as low as 0.016 seconds for 0.5-second samples—an improvement of 0.003 seconds. This makes it efficient for online steganalysis tasks, providing a superior balance between detection accuracy and efficiency under the constraint of short samples with low embedding rates.
@@
[label1]organization=LCPTS Laboratory, University of Sciences and Technology Houari Boumediene (USTHB),city=El-Alia, Bab-Ezzouar, postcode=16111, state=Algiers, country=Algeria
@@
[label2]organization=LSEA Laboratory, Department of Electrical Engineering, University of Medea,postcode=26000, state=Medea, country=Algeria
@@
[label3]organization=Cybersecurity Institute, School of Computer Science and Informatics, University of Liverpool,postcode=L69 3BX, state=Liverpool, country=UK
[…]
[label4]organization=Cyber Security Research Centre, School of Computing and Digital Media, London Metropolitan University,postcode=L69 3BX, state=London, country=UK
@@ 1 Introduction
Data hiding has garnered significant attention from researchers in recent years, spanning various disciplines. On the one hand, it serves to verify integrity and ensure information remains unaltered during transmission, as seen in watermarking [1, 2]. On the other hand, it can be utilized to conceal confidential information through unconventional means, exemplified by steganography [3, 4]. Steganography and steganalysis are two complementary aspects of covert communication. While steganography involves concealing secret information within apparently benign carriers such as images [5], text [6], and speech [3, 7, 8], steganalysis is focused on detecting and unveiling such concealed communication [9]. Over the years, the proliferation of Internet services has facilitated the transmission of multimedia data over digital networks. Among these services, voice over internet protocol (VoIP) has emerged as a prominent means of enabling real-time voice communication over the Internet. This increased reliance on VoIP, driven by the success of platforms like Skype, WhatsApp, and Zoom, has made it an attractive choice for concealing and transmitting hidden information [10]. The ubiquity and high-volume nature of VoIP traffic make it an appealing vector for steganographic purposes, presenting both opportunities for clandestine communication and challenges for cybersecurity professionals.
@@
In VoIP communications, speech is typically compressed using codecs, such as G.722, G.723, and G.729, to reduce bandwidth requirements. This compression process, which involves the quantization of speech parameters, inadvertently creates potential vulnerabilities that can be exploited for steganographic purposes. Malicious actors, including cybercriminals and extremist groups, may exploit these vulnerabilities by manipulating VoIP software to facilitate covert communications. This scenario presents a significant challenge to communication monitoring and network security. Steganography for VoIP can occur both in network protocol fields and payload data [11, 12]. However, while the former approach may provide some degree of concealment, it is the latter that takes precedence. This distinction arises from the fact that the practice of concealing information within network protocol fields, which encompass many layers of the open systems interconnection (OSI) model, ultimately provides less robust concealment capabilities. These protocol fields are typically public, and their data tends to remain fixed under most ordinary circumstances, rendering them less secure for covert communication [13]. On the contrary, the method of embedding data within payload data, characterized by its dynamic nature and temporal variations, offers a substantially higher degree of concealment. This dynamism makes it considerably more challenging to detect covert practices within VoIP streams, reinforcing its position as the preferred choice for steganographic endeavors [13, 14]. Additionally, as steganography methods become more sophisticated, particularly with the advent of AI-generated content, advanced steganalysis techniques are necessary to keep pace [15].
@@
An effective steganalysis in the context of VoIP streams should meet two crucial requirements [16, 17]. Firstly, it should operate in real-time, ensuring that the time required for detection is minimized due to the need for swift action against potential malicious activities. Secondly, the steganalysis method must be capable of detecting short samples of VoIP streams, as covert information may be hidden within brief segments of the communication. Adding to these two requirements, it should be sensitive enough to detect low embedding rates, as some covert data may be hidden with minimal changes to the host signal. Meeting these stringent requirements represents a significant and persistent challenge in developing practical and reliable VoIP steganalysis methods, which is essential to ensure their successful integration into real-world applications [16, 11]. In this work, we address the core research problem of designing a steganalysis approach for VoIP that can achieve real-time detection performance, maintain high accuracy on short speech segments, and remain effective under low embedding rates.
@@ 1.1 Motivation
With the rapid advancement of deep learning (DL), recent methods have leveraged deep models to automatically learn discriminative features. Various neural network architectures, primarily leveraging convolutional neural network (CNN)- and recurrent neural network (RNN)-based designs [18, 16, 14, 19], have been explored in VoIP steganalysis. These models are specifically designed to capture distinct codeword correlation features between cover and stego VoIP streams, achieving state-of-the-art (SOTA) detection outcomes. While effective in capturing sequential (RNNs) or local spatial (CNNs) patterns, these traditional deep learning architectures face inherent limitations when dealing with the relational structure of compressed VoIP data affected by steganography. QIM steganography, in particular, subtly alters the dependencies and correlations between line spectral frequency (LSF) codewords across frames and within their neighborhoods. CNNs, designed for grid-like data, struggle to explicitly model these non-Euclidean, inter-frame relationships. RNNs, while adept at sequences, might not efficiently capture the graph-like dependencies that extend beyond simple linear order or local windows, especially when the crucial information lies in the way codewords relate to each other rather than just their individual values or linear sequence.
Recently, significant attention has been directed toward adapting DL methods for graph-structured data, giving rise to the the emergence of graph neural networks (GNNs) as a prominent topic [20]. GNNs are uniquely suited for modeling complex relationships within non-Euclidean data by explicitly learning from graph structures, where nodes represent entities (e.g., LSF codeword frames) and edges represent their relationships (e.g., temporal dependencies). This intrinsic capability allows GNNs to capture both fine-grained local dependencies and high-level global patterns by aggregating information from connected neighborhoods. GNNs have demonstrated efficiency in representing and analyzing graph data across various computer vision fields, including action recognition [21], object tracking [22], and natural language processing, such as text classification [23]. However, there have been limited efforts in applying GNNs to steganalysis, primarily in the domains of images, as in [24, 25], and text, as in [26, 27]. This highlights a significant gap where GNNs’ relational modeling power could be particularly advantageous for VoIP steganalysis.
@@ 2 Related work
Vocoders, such as G.723 and G.729, are widely applied in VoIP communication with the aim of lowering decoding errors via the analysis-by-synthesis (AbS) approach, maintaining excellent speech quality while achieving high compression ratios. In contrast to conventional steganographic methods typically used for images or text, low bit-rate (LBR) VoIP streams, characterized by the AbS linear predictive coding technique, pose a unique challenge due to their minimal redundancy within the encoded speech. To address this, researchers have explored two distinct categories of data hiding techniques, each linked to different stages in the encoding process. The first category involves altering specific elements in the compressed speech stream, often utilizing methods like least significant bit (LSB) replacement, such as in [12]. However, this is complicated due to the limited data-hiding space in LBR speech. The second category focuses on embedding covert information during encoding, by altering encoding features, notably the linear predictive coding (LPC) [28], fixed codebook (FCB) [29], and adaptive codebook (ACB) [30].
@@
In the second category, QIM-based steganography is most commonly employed, especially within LPC domains. QIM steganography, originally developed by Chen et al. [31], is widely used in literature. It conceals steganograms by altering the quantization vector within the speech code, introducing minimal unperceptible distortion and offering high data hiding capabilities. Thus, when it comes to embedding information in VoIP streams, QIM emerges as a suitable scheme that poses a significant challenge for detection.
@@
On the one hand, research into QIM steganography detection methods has primarily concentrated on image carriers [32]. Subsequently, a limited number of research studies have been proposed in the field of steganalysis, particularly focusing on QIM-based steganography in VoIP, such as in [13, 33]. Due to the distinct characteristics of VoIP streams, the detection methods proposed for use in images cannot be directly applied within the VoIP context. Moreover, classical audio steganalysis methods based on statistical feature extraction in the uncompressed domain are inapplicable to VoIP. This is due to the minimal additional distortion introduced by VoIP QIM steganography in decoded speech signals, making it challenging to extract suitable features for steganalysis [34]. To address this issue, some research in the VoIP field has proposed the use of handcrafted statistical features to detect significant changes in codewords used in LBR speech codec streams caused by QIM steganography. For instance, Li et al. [35] examined the impact of QIM steganography on G.729, where it altered the quantization index of the LPC filter. They employed a statistical model to capture codeword distribution characteristics and, in conjunction with a support vector machine (SVM), formulated an efficient system for steganographic detection. In a related study, Li et al. [36] developed the index distribution characteristics (IDC) steganalysis method, employing codeword distribution histograms and first-order Markov chain-derived state transition probabilities as correlation features. These features, integrated with SVM, were utilized to classify steganographic samples with a focus on inter-frame transition probabilities. Expanding upon this concept, the studies by Li et al. [37] and Wu et al. [38] aim to detect steganograms embedded using QIM techniques. Both works employ principal component analysis (PCA) for reducing feature dimensionality and SVM for classification. Notably, Li et al. [37] focus on state transition probabilities of intra-frame codewords as features, while Wu et al. [38] suggested extracting features from the original speech signal by analyzing its distribution patterns and transmission likelihoods. Meanwhile, Yang et al. [39] employed temporal dependencies between codewords across frames to model a Codeword Bayesian Network (CBN), where the parameter learning process was guided by a Dirichlet prior. Yang et al. [40] proposed a fast steganalysis which involved mapping vector quantization codewords to a semantic space, followed by feature extraction using a hidden layer input into a softmax classifier. Such methods require meticulous feature design to align with the characteristics of speech data and specific information hiding techniques. Consequently, they struggle to effectively address generative steganography that produces high-quality stego speech. On the other hand, the proposed GraphSAGE-based algorithm addresses these challenges by utilizing relational data through GNNs and efficiently managing the minimal distortion and short sample lengths typical of VoIP streams.
@@
Besides, recent advancements in DL have led to improved robustness and effectiveness in steganalysis methods, which can be broadly categorized into two main categories: general steganalysis approaches and specific steganalysis approaches. In the first category, several methods have been introduced to detect a range of steganographic techniques. For example, Hu et al. [41] addressed heterogeneous parallel steganography (HPS) by introducing the steganalysis feature fusion network (SFFN). The architecture of SFFN comprises three core modules: a network for feature extraction, a component for feature integration, and a classification unit, effectively extracting and combining steganalysis features from HPS methods for dependable predictions. Building upon this, Wang et al. [42] proposed a fast and efficient steganalysis method which leveraged attention mechanisms. Their suggested method adeptly extracts key characteristics related to steganography exceptions and fuses these features targeting multiple steganography techniques utilized in HPS. Li et al. [43] introduced a unified detection framework, termed CBCA, which leverages codeword embedding along with bidirectional long short-term memory (Bi-LSTM) and CNN-based attention mechanisms. This approach is capable of identifying three distinct steganographic techniques concurrently. Wang et al. [44] proposed a steganalysis method aimed at identifying various steganographic schemes within compressed speech signals. Their framework introduces a codeword-distributed embedding component designed to generate compact representations from compressed codewords. To capture dependencies across different contexts, it employs two correlation extraction modules: a global-guided unit that incorporates Bi-LSTM and multi-head self-attention, and a local-guided unit built using the convolutional block Attention module (CBAM), enabling the modeling of both global and local correlations before and after data hiding. Detection is accomplished through fully connected layers, ensuring accurate identification. Li et al. [15] introduced a generic frame-level steganalysis approach for LBR steganography, utilizing dual-domain feature representation and a Transformer-based model named Stegaformer. The proposed method consists of two modules, dual-domain representation and Stegaformer. Li et al. [45] introduced SANet, an independent steganalysis network for speech encoding and steganography. This approach unifies compressed speech outputs from various codecs by converting them into a common uncompressed-domain format, introducing an intermediate representation. A Bi-LSTM neural network is developed to extract steganography-sensitive characteristics through collaborative correlation features. This approach achieves SOTA detection performance for various steganography algorithms across different speech encoders. Recently, Wang et al. [46] proposed E-SWAN, a deep learning-based sliding window analysis network that combines LSTM and convolutional modules for real-time VoIP speech steganalysis. The model demonstrates competitive performance in detecting two steganography methods. Similarly, Lin et al. [47] introduced a steganalysis approach that integrates Bi-LSTM, 3D convolution and attention mechanisms to detect two QIM-based steganography variants in VoIP streams. Although these algorithms offer some capability for broad steganalysis applications, their performance often falters when confronted with specific steganographic techniques, particularly QIM. In such specialized scenarios, these general methods typically exhibit lower accuracy, highlighting the need for more targeted approaches.
@@
Consequently, several researchers prefer to exploit the power of DL to design specialized steganalysis algorithms, particularly geared toward QIM-based steganography. For example, Lin et al. [16] identified four robust codeword correlation patterns and introduced a steganalysis model based on RNN for improved detection accuracy. Nonetheless, their model, RNN-SM, centred on global contextual information, utilizing a two-layer LSTM. Yang et al. [34] proposed a novel approach to enhance local representations that combined RNN and CNN, resulting in the CNN-LSTM model, which outperformed RNN-SM in the detection of QIM-based steganography. Yang et al. [13] incorporated an attention mechanism within a hierarchical convolutional structure, further enhancing steganalysis results. In a subsequent study [48], the same authors presented a lightweight neural architecture called the fast correlation extract model (FCEM), which integrates positional encoding and multi-head attention to capture correlation features. Qiu et al. [18] developed an efficient steganalysis that incorporates a codeword embedding layer to capture dense representations, employs a bidirectional LSTM layer and incorporates a gated attention module to model contextual feature distributions. Yang et al. [14] designed a multi-channel convolutional sliding window mechanism to capture inter-frame dependencies between a target frame and its adjacent frames. Recently, the utilization of a transformer encoder [49, 50, 51] and federated learning [52, 53] has been implemented for the detection of QIM-based steganography in G.729 speech encoders. These two approaches demonstrated SOTA detection performance. Zhang et al. [54] tackled the challenge of identifying payload locations in QIM-based VoIP steganography by proposing an LSTM-based steganalysis method named SPM. To provide a structured comparison and highlight the evolution of these key approaches, Table 1 presents a summary of the works most relevant to our research, detailing their core techniques and limitations.
Table 1: A comparative summary of key SOTA methods in QIM-based VoIP steganalysis, representing the works most closely related to our proposed approach.
Ref. Method / Model Key contribution Limitations addressed by our work
Li et al. [35] Statistical model + SVM Captures distribution characteristics of LPC quantization indices in G.729 Requires handcrafted features; limited robustness to low embedding rates
Li et al. [36] Statistical Features + SVM Used codeword histograms and Markov probabilities to capture inter-frame correlations. Relies on manually designed features; limited robustness to low embedding rates
Li et al. [37] Statistical Features + PCA + SVM Focused on state transition probabilities of intra-frame codewords. Requires meticulous feature engineering and is less adaptable than end-to-end models; limited robustness to low embedding rates
Wu et al. [38] Statistical Features + PCA + SVM Extracted features from the original speech signal’s distribution patterns. Feature design is specific to known signal characteristics; Computationally expensive; weaker for real-time detection.
Yang et al. [39] Codeword bayesian network (CBN) Modeled temporal dependencies across frames using a probabilistic graphical model. Relies on a specific statistical model (CBN) that is less flexible than learned representations.
Yang et al. [40] Semantic mapping + softmax classifier Projects VQ codewords to semantic space for classification Lightweight but less accurate for subtle embeddings
Lin et al. [16] RNN-SM (Two-layer LSTM) Captures global contextual correlations Models data purely sequentially; less effective at capturing complex, non-local dependencies.
Yang et al. [34] CNN + LSTM Combines local CNN with temporal RNN features Grid-based, CNN and sequential models, do not explicitly capture the complex relational structure of VoIP data; Higher complexity
Yang et al. [13] Hierarchical CNN + Attention Enhanced detection using multi-level attention Limited by the inherently local receptive field of CNNs, which struggles to capture complex, long-range dependencies
Yang et al. [48] FCEM (Lightweight CNN + Attention) Efficient correlation extraction with positional encoding Attention works on sequences; weaker in capturing structural relationships
Qiu et al. [18] Bi-LSTM + gated Attention Dense contextual feature representation Sequential model; lacks relational aggregation over graph structures
Zhang et al. [54] LSTM-based SPM Identifies payload locations in VoIP Task-specific; not generalized for broad steganalysis
Yang et al. [14] Multi-channel sliding CNN Window Captures local inter-frame dependencies between adjacent frames Sliding window has fixed local receptive field; cannot capture longer-range dependencies
Zhang et al. [49] Transformer encoder Leveraged the power of self-attention from Transformers for steganalysis. May not inherently model the direct temporal adjacency in speech frames as effectively as a graph structure.
Tian et al. [52] Federated learning Introduced a secure, collaborative learning framework for steganalysis. Focuses on the training paradigm rather than the core model architecture for feature extraction.
Our Work GNN First to model VoIP streams as graphs to capture hierarchical relational data. Explicitly models non-Euclidean, relational dependencies between speech frames, overcoming the limitations of both manual feature engineering and simpler sequential/grid-based models.
@@
The mentioned studies underscore the adoption of DL-based methods as the benchmark for real-time VoIP steganalysis, demonstrating superior detection capabilities and efficiency in comparison to handcrafted features. The integration of DL with graph data has led to the widespread adoption of GNNs. These DL-based models, specialized for graph domains, are crafted to capture representations of graph data adeptly, proficiently handling tasks centred on both individual nodes and entire graphs. GNNs capture relationships within graphs by exchanging messages among nodes and maintaining a state that encodes information from any connected neighborhood through edges [55]. Furthermore, GNNs exhibit an enhanced capability to leverage global information, contributing to their compelling performance. While GNNs have been recently employed in various studies for image [24, 25] and text steganalysis [26, 27, 56], there is a noticeable gap in research focusing on steganalysis in the context of VoIP.